OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

put it into sloper-intercept form y = mx + b

OpenStudy (anonymous):

y = mx + b 2x + 5y = -10 5y = -2x -10 whats next?

OpenStudy (anonymous):

im not sure.. divide -2?

OpenStudy (anonymous):

nope you want y alone not x

OpenStudy (anonymous):

so divide 5

OpenStudy (anonymous):

yes what do you get?

OpenStudy (anonymous):

y=-2x-5?

OpenStudy (anonymous):

nope try again

OpenStudy (anonymous):

what is 10/5?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

oops y=2x-2

OpenStudy (anonymous):

ok -2/5?

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

ok what is the equation then

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

*face palm*

OpenStudy (anonymous):

5y = -2x -10 divide everything by 5 what you get

OpenStudy (anonymous):

y=-0.4-2

OpenStudy (anonymous):

whatever forget it

OpenStudy (anonymous):

lol why?

OpenStudy (anonymous):

why did x run away?

OpenStudy (anonymous):

^^^ LOLOL